OpenStudy (anonymous):

Plz help. An answer and an explanation would be Great! Thanks Find the 5th terms of (y - 4)8. -17,920y4 17,020y4 17,920y4 -17,020y4

OpenStudy (anonymous):

\[\left( y-4 \right)^8\] \[T _{r+1}=C _{r}^{8}*y ^{8-r}\left( -4 \right)^r\] put r+1=5,r=5-1=4 \[C_{4}^{8}=\frac{ 8*7*6*5 }{ 4*3*2*1 }=?\]

OpenStudy (anonymous):

The question mark equals 151.6666667, but that is not listed as an answer. Is there another step?

OpenStudy (anonymous):

\[T _{5}=C _{4}^{8}y ^{8-4}\left( -1 \right)^4*4^4\] \[=70 y^4*256=?\]

OpenStudy (anonymous):

I really have no clue what to do. Can you explain a bit and give the answer so I can work backwards? Thanks

OpenStudy (anonymous):

\[=17920 y^4 \]

OpenStudy (anonymous):

\[70 y^4 *256=17920 y^4 \] i have used binomial theorem for any positive integral index.